Solve the following equations using appropriate substitutions

a) y' = 5y +( e^−3x)(y^−3)


b) xdy = y(ln x − ln y)dx, y(1) = 4, x > 0.



1
Expert's answer
2021-05-07T09:56:00-0400

Solution.

part a)

"y'-5y-e^{-3x}y^{-3}=0"

Rewrite the

original differential equation as follows:

"-5y^4e^{3x}+y^3y'e^{3x}=1"

This is the Bernoulli equation for n = 4.

Substitution "z =y^4."

Then: "z'=4y^3y',"

and therefore the equation is rewritten as

"-5ze^{3x}+\frac{1}{4}z'e^{3x}=1."

We solve this equation by the method of variation of an arbitrary constant.

This is a non-homogeneous equation. Consider the corresponding homogeneous equation:

"-5ze^{3x}+\frac{1}{4}z'e^{3x}=0."

Solving it, we get:

"z'=20z."

Integrating, we get:

"\ln{z}=20x,\newline
z=e^{20x}."

We are now looking for a solution of the original equation in the form:

"z(x)=C(x)e^{20x},\newline
 z '(x) = C (x) e ^{20 x} + C (x) (e ^{20 x} ) '."

So,

"C'(x)= 4 e ^{-23 x}"

Integrating, we obtain:

"C(x)=-\frac{4}{23}e^{-23x}."

From the condition

"z(x)=-\frac{4}{23}e^{-23x}e^{20x}."

Since "z=y^4,"  we get:

"y^4=Ce^{20x}-\frac{4}{23}e^{-3x}."

part b)


"xy′=y(\ln x−\ln y)."

We make substitution



"u(x)=\frac{y(x)}{x}."

Then "y(x)=u(x)x" and

"y'=u'x+u."

We will have

"x^2u'+xu=ux(\ln x-\ln (ux)),"

or "x^2u'+xu\ln u+xu=0."

Solve this equation:

"\frac{u'}{(\ln u+1)u}=-\frac{1}{x}," or

"\frac{du}{(\ln u+1)u}=-\frac{dx}{x}."

"\int\frac{du}{(\ln u+1)u}=-\int\frac{dx}{x},"

"\ln ({\ln u +1})=-\ln x +C,"

where C is some constant.

From here "u = e^{\frac{C_1}{x}-1}."

And the solution of given equation is


"y = xe^{\frac{C_1}{x}-1}."

y(1)=4, then

"e^{C_1-1}=4,\newline
C_1-1=\ln 4,\newline
C_1=\ln 4+1,\newline
y=xe^{\frac{\ln 4 +1}{x}-1}."